Maleic acid is an organic compound composed of 41.39% C, 3.47% H, and the rest oxygen. If 0.129 mol of maleic acid has a mass of 15.0 g, what are the empirical and molecular formulas of maleic acid?


1
Expert's answer
2021-10-05T23:58:42-0400

Number of moles of maleic acid"=0.168mol =\\frac{19.5g}{M\\frac{g}{mol}}\n\n\u200b"


So molar mass of maleic acid"=M=\frac{19.5}{0.168}=116.07\frac{g}{mol}"


Let atoms of CC be xx


Let atoms of HH be yy


Let atoms of OO be zz


"12\times x=0.4139\times116.07"


"1\times y=0.0347\times 116.07"

"16\times z=0.5514\times116.07"


"x\approx4x≈4"


"y\approx4y≈4"


"z\approx4z≈4"


Hence molecular formula of maleic acid is "C_4H_4O_4"


And the empirical formula is CHOCHO